Q4) Sellstuff Ltd. is valued at £18 millions and currently earns a 15% return on capital. Losses from pure risks at Sellstuff are usually around £180,000. For the forthcoming year Sellstuffs insurers have quoted an insurance premium of £700,000 to cover all losses. REQUIRED: (making any necessary assumptions) a) Calculate Sellstuffs opportunity cost of insurance (8 marks) b) Calculate the opportunity cost if management at Sellstuff set up a E2,400,000 fund to cover potential losses. The fund would be invested at a risk free 5% rate of interest and cost £300,000 to administer. (12 marks)
On the basis of the information and your calculations so far for this question, offer your opinions on whether the managers at Sellstuff should choose the full insurance (as in a)) or the fund (as in b)). (10 marks) c) Explain what information OTHER than that provided is relevant to the situation and how it might influence the decision to insure or self-insure. (10 marks) d) Suggest and explain details of alternative ways that Sellstuffs pure loss exposure could be handled that might be more sensible than either full insurance or full self-insurance (10 marks). e)
Answer #1

a) total insurance premium: 700000

return on capital: 15%

If this amount is not used for paying premium for insurance, We can assume that this will be invested in the business itself which gives 15% return.

Return from investing in own business: 700000*0.15 = 105000

Thus the opportunity cost of insurance is 105000 euros.

b) Now we are setting up a fund of 2400000 euros which will be used to pay up for the losses. But here we will gain a 5% return unlike the previous case. Also it entails an expenditure of 300000 euros for administering the fund on which we wont earn any return.

Thus the total opportunity cost would be the sum of 1) opportunity cost of fund i.e 24 00000 euros and 2) opportunity cost of adimistration i.e 300000 .

For the fund the opportunity return would be 10% ( 15-5)

And for the administration fund the opportunity cost would be 15%.

Now Opportunity cost of the fund= 2400000 * 0.1= 240000

opportunity cost of the administered fund = 300000*0.15= 45000

total= 240000+45000 = 285000

c) based on the calculations as we saw before, in case where the company opts for the insurance, the opportunity cost is 105000 euros whereas when the company sets up the fund the opportunity cost is 285000. Now we clearly see that in the insurance case the company's opportunity cost is lower. If the company selects this method, it can achieve its objective of covering the losses with minimum investment. Thus the insurance option is preferable.

d) if the probability of losses were given, we could figure out the expected losses for the next year and determine a right premium for the losses. In the absence of any probability data, insurance seems better.

e) One option that can be selected is diversifying the business and investing into other business whose returns are negatively related to the pure risks of the company. So, In case it suffers losses from Sellstuff, it might still gain from the other diversified businesses and in total may stay more profitable.